The kinetic energy of a charged particle decreases by 10J as it moves frm a pt. at  potential 100 V to 200 V .Find charge on particle.

Grade:Upto college level

1 Answers

vikas askiitian expert
13 years ago

electric potential energy=q.V

  applying energy conservation between two points

  V1q + KE1 =V2q +KE2

           KE1-KE2=q(V2-V1)

          q=10/100=0.1c